# An additive two-level parallel variant of the DMRG algorithm with coarse-space correction

### v1

@@ 3.1. The Second-Level Minimization in the A2DMRG Algorithm 2
An essential feature of A2DMRG Algorithm 2 is the solution of a minimization problem of dimension at most $d\in\mathbb{N}$ (recall that $d$ is the order of the underlying tensor space). This minimization problem yields an optimal linear combination of the $d$ tensor train decompositions arising from the first-level parallel one-site/two-site DMRG micro-iterations. We now discuss the resolution of this second-level minimization problem for Rayleigh quotient-based energy functionals $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{R}$ of the form (2), i.e.,
@@
$\forall\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{U})=\frac{\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle}{\|\mathbb{U}\|^{2}},$
@@
where $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ is a symmetric operator with a simple lowest eigenvalue. In this case, the second-level minimization problem (13) for the $n^{\rm th}$ global iteration takes the form
@@
(14) $\underset{0\neq\mathbb{U}\in\mathcal{Y}_{n}}{\min}\;\frac{\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle}{\|\mathbb{U}\|^{2}}\qquad\text{where }~{}\mathcal{Y}_{n}:=\text{span}\Big{\{}\tau\big{(}\widetilde{\mathbb{Y}}_{i}^{(n+1)}\big{)}\colon\quad i\in\{1,\ldots,d\}\Big{\}}.$
@@
Using the fact that the tensors $\{\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})\}_{i=1}^{d}$ span the space $Y_{n}$ by definition, we can express the minimization problem (14) as a matrix problem of the form
@@
(15) $\underset{\begin{subarray}{c}\mathbb{c}=(c_{1},\ldots,c_{d})\in\mathbb{R}^{d}\\[5.0pt] {\sum_{i=1}^{d}c_{i}\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})}\neq 0\end{subarray}}{\min}\;\frac{\langle\mathbb{c},\widehat{\mathbb{A}}\mathbb{c}\rangle}{\langle\mathbb{c},\widehat{\mathbb{S}}\mathbb{c}\rangle}.$
Here, the symmetric matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d\times d}$ and the overlap matrix $\widehat{\mathbb{S}}\in\mathbb{R}^{d\times d}$ are given by
@@
(16) $\forall i,j\in\{1,\ldots,d\}\colon\qquad[\widehat{\mathbb{A}}]_{i,j}$ $:=\langle\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)}),\mathbb{A}\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})\rangle$
@@
$\forall i,j\in\{1,\ldots,d\}\colon\qquad[\widehat{\mathbb{S}}]_{i,j}$ $:=\langle\tau(\widetilde{\mathbb{Y}}_{j}^{(n+1)}),\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})\rangle.$
@@
Since $\widehat{\mathbb{S}}$ is a symmetric matrix, we can introduce the eigendecomposition $\widehat{\mathbb{S}}=\widehat{\mathbb{U}}_{\mathbb{S}}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}$ and deduce that
@@
$\underset{\begin{subarray}{c}\mathbb{c}=(c_{1},\ldots,c_{d})\in\mathbb{R}^{d}\\[5.0pt] {\sum_{i=1}^{d}c_{i}\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})}\neq 0\end{subarray}}{\min}\;\frac{\langle\mathbb{c},\widehat{\mathbb{A}}\mathbb{c}\rangle}{\langle\mathbb{c},\widehat{\mathbb{S}}\mathbb{c}\rangle}$ $=\underset{\begin{subarray}{c}\mathbb{c}=(c_{1},\ldots,c_{d})\in\mathbb{R}^{d}\\[5.0pt] {\sum_{i=1}^{d}c_{i}\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})}\neq 0\end{subarray}}{\min}\;\frac{\langle\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}{\langle\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}.$
@@
$\underset{\begin{subarray}{c}\mathbb{c}=(c_{1},\ldots,c_{d})\in\mathbb{R}^{d}\\[5.0pt] {\sum_{i=1}^{d}c_{i}\tau(\widetilde{\mathbb{Y}}_{i}^{(n+1)})}\neq 0\end{subarray}}{\min}\;\frac{\langle\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}{\langle\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}$ $=\underset{0\neq\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\in\mathbb{R}^{d}}{\min}\;\frac{\langle\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}{\langle\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}$
$=\underset{0\neq\widehat{\mathbb{c}}\in\mathbb{R}^{d}}{\min}\;\frac{\langle\widehat{\mathbb{c}},\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}\widehat{\mathbb{c}}\rangle}{\langle\widehat{\mathbb{c}},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{c}}\rangle},$
@@
$\big{(}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{-1/2}\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{-1/2}\big{)}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{1/2}\widehat{\mathbb{c}}=\lambda\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{1/2}\widehat{\mathbb{c}}.$
Thus, the sought-after minimizer of the second-level minimization problem (14) is given by $\sum_{i=1}^{d}c_{i}^{*}\tau\big{(}\widetilde{\mathbb{Y}}_{i}^{(n+1)}\big{)}$ with the vector $\mathbb{c}^{*}=(c_{1},\ldots,c_{d})$ being obtained by computing the lowest eigenpair of a generalized eigenvalue problem involving the matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d\times d}$ and the eigendecomposition factors of the overlap matrix $\widehat{\mathbb{S}}$.
If, on the other hand, the diagonal matrix $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ is rank-deficient, then we can decompose $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ and $\widehat{\mathbb{U}}_{\mathbb{S}}$ as
$\widehat{\mathbb{\Sigma}}_{\mathbb{S}}=\begin{bmatrix}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}&0\\ 0&0\end{bmatrix},\hskip 5.69054pt\widehat{\mathbb{U}}_{\mathbb{S}}=\begin{bmatrix}\widehat{\mathbb{U}}_{\mathbb{S}}^{+}&\widehat{\mathbb{U}}_{\mathbb{S}}^{0}\end{bmatrix}~{}\text{ where }\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}\in\mathbb{R}^{p\times p}\text{ is full-rank and }\widehat{\mathbb{U}}_{\mathbb{S}}^{+}\in\mathbb{R}^{d\times p}.$
Using now the matrix $\widehat{\mathbb{U}}_{\mathbb{S}}^{+}\in\mathbb{R}^{d\times p}$, we can construct the following $p$-dimensional orthogonal basis for the approximation space $Y_{n}$:
$\left\{\sum_{j=1}^{d}[\widehat{\mathbb{U}}_{\mathbb{S}}^{+}]_{jk}\tau(\widetilde{\mathbb{Y}}^{(n+1)}_{j})\colon\quad k\in\{1,\ldots,p\}\right\}.$
Proceeding as before, we deduce that the second-level minimization problem (14) can be expressed in the form
(17) $\underset{0\neq\mathbb{U}\in Y_{n}}{\min}\;\frac{\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle}{\|\mathbb{U}\|}=\underset{\begin{subarray}{c}\mathbb{b}=(b_{1},\ldots,b_{p})\in\mathbb{R}^{p}\\ \mathbb{b}\neq 0\end{subarray}}{\min}\;\frac{\langle\mathbb{b},(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}^{+}\mathbb{b}\rangle}{\langle\mathbb{b},\mathbb{\Sigma}_{\mathbb{S}}^{+}\mathbb{b}\rangle}.$
@@
(18) $\big{(}(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}(\widehat{\mathbb{U}}_{\mathbb{S}}^{+})^{*}\widehat{\mathbb{A}}\widehat{\mathbb{U}}_{\mathbb{S}}(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}\big{)}(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{1/2}\widehat{\mathbb{b}}=\lambda(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{1/2}\widehat{\mathbb{b}}.$
Consequently, the sought-after minimizer of the second-level problem (14) is given by
$\sum_{\ell=k}^{p}\widehat{b}^{*}_{k}\sum_{j=1}^{d}[\widehat{\mathbb{U}}_{\mathbb{S}}^{+}]_{jk}\widetilde{\mathbb{Y}}^{(n+1)}_{j},$
with the vector $\widehat{\mathbb{b}}^{*}=(b^{*}_{1},\ldots,b^{*}_{p})$ being obtained once again by computing the lowest eigenpair of an eigenvalue problem involving the matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d\times d}$ and the eigendecomposition factors of the overlap matrix $\widehat{\mathbb{S}}$. We remark that in the case when the diagonal matrix $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ is ill-conditioned but not rank-deficient, we can similarly decompose $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ and $\widehat{\mathbb{U}}_{\mathbb{S}}$ using a tolerance parameter $\epsilon>0$.
@@
(19) $\forall\mathbb{U}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\mathcal{J}(\mathbb{U})=\frac{1}{2}\langle\mathbb{U},\mathbb{A}\mathbb{U}\rangle-\langle\mathbb{F},\mathbb{U}\rangle,$
with $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ symmetric, positive and $\mathbb{F}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$.

### v2

@@ 3.1. The Second-Level Minimization in the A2DMRG Algorithm 2
An essential feature of the A2DMRG Algorithm 2 is the solution of a minimization problem of dimension at most $d+1\in\mathbb{N}$ for the one-site case and $d\in\mathbb{N}$ for the two-site case (recall that $d$ is the order of the underlying tensor space). This minimization problem yields an optimal linear combination of the $d$ (respectively, d1d-1) tensor train decompositions arising from the first-level one-site (respectively, two-site) DMRG micro-iterations together with the previous global iterate. We now discuss the resolution of this second-level minimization problem for Rayleigh quotient-based energy functionals $\mathcal{J}\colon U\subset\mathbb{R}^{n_{1}\times\ldots\times n_{d}}\rightarrow\mathbb{R}$ of the form (2), i.e.,
@@
$\forall\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\setminus\{0\}\colon\qquad\mathcal{J}(\mathbb{X})=\frac{\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle}{\|\mathbb{X}\|^{2}},$
@@
where $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ is a symmetric operator with a simple lowest eigenvalue. For simplicity, we assume we are in the one-site setting (the two-site being identical up to a modification of indices). In this case, the second-level minimization problem (13) for the $n^{\rm th}$ global iteration takes the form
@@
(14) $\underset{0\neq\mathbb{X}\in\mathcal{Y}_{n}}{\min}\;\frac{\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle}{\|\mathbb{X}\|^{2}}\qquad\text{where }~\mathcal{Y}_{n}:=\text{span}\Big\{\tau\big(\widetilde{\mathbb{U}}^{(n+1),i}\big)\colon\quad i\in\{0,\ldots,d\}\Big\}.$
@@
Using the fact that the tensors $\{\tau(\widetilde{\mathbb{U}}^{(n+1),i})\}_{i=0}^{d}$ span the space $\mathcal{Y}_{n}$ by definition, we can express the minimization problem (14) as a matrix problem of the form
@@
(15) $\underset{\begin{subarray}{c}\mathbb{c}=(c_{0},\ldots,c_{d})\in\mathbb{R}^{d+1}\\[5.0pt] {\sum_{i=0}^{d}c_{i}\tau(\widetilde{\mathbb{U}}^{(n+1),i})}\neq 0\end{subarray}}{\min}\;\frac{\langle\mathbb{c},\widehat{\mathbb{A}}\mathbb{c}\rangle}{\langle\mathbb{c},\widehat{\mathbb{S}}\mathbb{c}\rangle}.$
Here, the symmetric matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d+1\times d+1}$ and overlap matrix $\widehat{\mathbb{S}}\in\mathbb{R}^{d+1\times d+1}$ are
@@
(16) $\forall i,j\in\{1,\ldots,d+1\}\colon\qquad[\widehat{\mathbb{A}}]_{i,j}$ $:=\big\langle\tau(\widetilde{\mathbb{U}}^{(n+1),{i-1}}),\mathbb{A}\;\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})\big\rangle$
@@
$\forall i,j\in\{1,\ldots,d+1\}\colon\qquad[\widehat{\mathbb{S}}]_{i,j}$ $:=\big\langle\tau(\widetilde{\mathbb{U}}^{(n+1),{i-1}}),\tau(\widetilde{\mathbb{U}}^{(n+1),{j-1}})\big\rangle.$
@@
Since $\widehat{\mathbb{S}}$ is symmetric, we can introduce the eigendecomposition $\widehat{\mathbb{S}}=\widehat{\mathbb{V}}_{\mathbb{S}}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}$ and deduce that
@@
$\underset{\begin{subarray}{c}\mathbb{c}=(c_{0},\ldots,c_{d})\in\mathbb{R}^{d+1}\\[5.0pt] {\sum_{i=0}^{d}c_{i}\tau(\widetilde{\mathbb{U}}^{(n+1),i})}\neq 0\end{subarray}}{\min}\;\frac{\langle\mathbb{c},\widehat{\mathbb{A}}\mathbb{c}\rangle}{\langle\mathbb{c},\widehat{\mathbb{S}}\mathbb{c}\rangle}$ $=\underset{\begin{subarray}{c}\mathbb{c}=(c_{0},\ldots,c_{d})\in\mathbb{R}^{d+1}\\[5.0pt] {\sum_{i=0}^{d}c_{i}\tau(\widetilde{\mathbb{U}}^{(n+1),i})}\neq 0\end{subarray}}{\min}\;\frac{\langle\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}{\langle\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}.$
@@
$\underset{\begin{subarray}{c}\mathbb{c}=(c_{0},\ldots,c_{d})\in\mathbb{R}^{d+1}\\[5.0pt] {\sum_{i=0}^{d}c_{i}\tau(\widetilde{\mathbb{U}}^{(n+1),i})}\neq 0\end{subarray}}{\min}\;\frac{\langle\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{U}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}{\langle\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}$ $=\underset{0\neq\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\in\mathbb{R}^{d+1}}{\min}\;\frac{\langle\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}{\langle\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\mathbb{c}\rangle}$
$=\underset{0\neq\widehat{\mathbb{c}}\in\mathbb{R}^{d+1}}{\min}\;\frac{\langle\widehat{\mathbb{c}},\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}\widehat{\mathbb{c}}\rangle}{\langle\widehat{\mathbb{c}},\widehat{\mathbb{\Sigma}}_{\mathbb{S}}\widehat{\mathbb{c}}\rangle}$,
@@
$\big(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{-1/2}\widehat{\mathbb{V}}_{\mathbb{S}}^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{-1/2}\big)\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{1/2}\widehat{\mathbb{c}}=\lambda\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{1/2}\widehat{\mathbb{c}}$.
Thus, the minimizer of the second-level minimization problem (14) is given by $\sum_{i=0}^{d}c_{i}^{*}\tau\big(\widetilde{\mathbb{U}}^{(n+1),i}\big)$ with the vector $\mathbb{c}^{*}=(c_{0}^{*},\ldots,c_{d}^{*})$ being obtained by computing the lowest eigenpair of a generalized eigenvalue problem involving the matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d+1\times d+1}$ and the eigendecomposition factors of the overlap matrix $\widehat{\mathbb{S}}$.
If, on the other hand, the diagonal matrix $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ is rank-deficient, say, of rank $p<d+1$, then we can decompose $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ and $\widehat{\mathbb{U}}_{\mathbb{S}}$ as
$\widehat{\mathbb{\Sigma}}_{\mathbb{S}}=\begin{bmatrix}\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}&0\\ 0&0\end{bmatrix},\hskip 4.2679pt\widehat{\mathbb{U}}_{\mathbb{S}}=\begin{bmatrix}\widehat{\mathbb{V}}_{\mathbb{S}}^{+}&\widehat{\mathbb{V}}_{\mathbb{S}}^{0}\end{bmatrix}~\text{ with }\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+}\in\mathbb{R}^{p\times p}\text{ full-rank and }\widehat{\mathbb{V}}_{\mathbb{S}}^{+}\in\mathbb{R}^{d+1\times p}.$
Using now the matrix $\widehat{\mathbb{V}}_{\mathbb{S}}^{+}\in\mathbb{R}^{d+1\times p}$, we can construct the following $p$-dimensional orthogonal basis for the approximation space $\mathcal{Y}_{n}$:
$\left\{\sum_{j=0}^{d}[\widehat{\mathbb{V}}_{\mathbb{S}}^{+}]_{j+1,k}\;\tau(\widetilde{\mathbb{U}}^{(n+1),j})\colon\quad k\in\{1,\ldots,p\}\right\}.$
Consequently, the second-level minimization problem (14) can be expressed as
(17) $\underset{0\neq\mathbb{X}\in\mathcal{Y}_{n}}{\min}\;\frac{\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle}{\|\mathbb{X}\|}=\underset{\begin{subarray}{c}\mathbb{b}=(b_{1},\ldots,b_{p})\in\mathbb{R}^{p}\\ \mathbb{b}\neq 0\end{subarray}}{\min}\;\frac{\langle\mathbb{b},(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}^{+}\mathbb{b}\rangle}{\langle\mathbb{b},\mathbb{\Sigma}_{\mathbb{S}}^{+}\mathbb{b}\rangle}.$
@@
(18) $\big((\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}(\widehat{\mathbb{V}}_{\mathbb{S}}^{+})^{*}\widehat{\mathbb{A}}\widehat{\mathbb{V}}_{\mathbb{S}}(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{-1/2}\big)(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{1/2}\widehat{\mathbb{b}}=\lambda(\widehat{\mathbb{\Sigma}}_{\mathbb{S}}^{+})^{1/2}\widehat{\mathbb{b}}$.
Thus, the sought minimizer of the second-level minimization problem (14) is given by
$\sum_{k=1}^{p}\widehat{b}^{*}_{k}\sum_{j=0}^{d}[\widehat{\mathbb{V}}_{\mathbb{S}}^{+}]_{j+1,k}\;\tau(\widetilde{\mathbb{Y}}^{(n+1),j}),$
with the vector $\widehat{\mathbb{b}}^{*}=(b^{*}_{1},\ldots,b^{*}_{p})$ being obtained once again by computing the lowest eigenpair of an eigenvalue problem involving the matrix $\widehat{\mathbb{A}}\in\mathbb{R}^{d+1\times d+1}$ and the eigendecomposition factors of the overlap matrix $\widehat{\mathbb{S}}$. We remark that in the case when the diagonal matrix $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ is ill-conditioned but not rank-deficient, we can similarly decompose $\widehat{\mathbb{\Sigma}}_{\mathbb{S}}$ and $\widehat{\mathbb{U}}_{\mathbb{S}}$ using a tolerance parameter $\epsilon>0$.
@@
(19) $\forall\mathbb{X}\in U=\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\colon\qquad\mathcal{J}(\mathbb{X})=\frac{1}{2}\langle\mathbb{X},\mathbb{A}\mathbb{X}\rangle-\langle\mathbb{F},\mathbb{X}\rangle,$
with $\mathbb{A}\colon\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}\rightarrow\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$ symmetric and positive and $\mathbb{F}\in\mathbb{R}^{n_{1}\times n_{2}\times\ldots n_{d}}$. In this case, the local updates of A2DMRG (Step 2 in Algorithm 2) will consist of solving linear systems involving an orthogonal projection of $\mathbb{A}$ (see, e.g., [19] for details), while the second level minimization (Step 3 in Algorithm 2) will consist of the solution to a linear system involving the projection of $\mathbb{A}$ on the coarse $d+1$ or $d$-dimensional space spanned by the local updates.